# B-cos LM: Efficiently Transforming Pre-trained Language Models for Improved Explainability

### v3

@@ 4.2 Automatic Evaluation Results
Figure 2: Mean accuracy of conventionally fine-tuned, Saloss and B-cos BERT models averaged over three runs. We use B=1.5, 1.25, and 1.5 for AG News, IMDB, and HateXplain, respectively. B-cos models perform comparably to conventional models on most tasks.
@@
Table 2 shows the faithfulness scores for post-hoc explanation methods on conventionally fine-tuned and Saloss BERT models, as well as B-cos explanations from B-cos BERT. The results show that B-cos explanations are consistently and substantially more faithful than post-hoc methods across all models and datasets. On average, B-cos explanations outperform the strongest post-hoc methods on conventional models by 14.63 points in Comp and achieve negative Suff scores, indicating that the identified important tokens alone enable even more confident predictions. B-cos also shows significant gains in SeqPG. While Saloss improves faithfulness for some post-hoc methods over conventional models, it still underperforms compared to B-cos LMs by a large margin. Similar trends are observed for DistilBERT and RoBERTa (Appendix H) as well, further strengthening our findings. Although we do not include rationale-based models in the main experiments because they typically require additional supervision, a supplementary comparison in Appendix I shows that B-cos BERT still outperforms a rationale-based model on HateXplain.
@@
Table 2: Faithfulness evaluation for conventionally fine-tuned BERT, Saloss BERT and B-cos BERT across three datasets. We use B=1.5, 1.25, and 1.5 for AG News, IMDB and HateXplain, respectively. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
@@ 4.3 Human Evaluation
Contrary to previous B-cos studies that rely solely on automatic evaluations to assess explanations, we conduct the first human study to better evaluate the human interpretability and agreement of B-cos explanations. We compare B-cos explanations against three strong post-hoc explanation methods on the conventional BERT model.
Following the practice in Enguehard (2023) and Yue et al. (2022), we randomly select 50 instances, respectively, from AG News and HateXplain where the B-cos and conventional models make the same prediction. Five annotators then rate the explanations in terms of human interpretability (how well they understand them) and human agreement (how much they agree with them) on a scale of 1-5. Before the annotation began, the annotators were provided with a clear description of the evaluation task, metrics, and rating scales. They were also shown example annotations along with the reasoning behind each rating to help them better understand the evaluation criteria. Further details on the evaluation criteria, rating scales, annotator instructions, and example annotations can be found in Appendix J.
@@
Figure 3 shows that B-cos explanations have a better human interpretability and exhibit greater alignment with human reasoning than post-hoc methods, even though they are not directly optimized for human agreement. Paired t-tests with a Bonferroni-corrected significance level $\alpha=\frac{0.05}{6}=0.008\overline{3}$ (Bonferroni, 1936) shows that the improvements of B-cos explanations are statistically significant ($p<\alpha$) for both metrics.
@@ 4.4 Qualitative Analysis
Figure 4 provides an example of B-cos and other (post-hoc) explanations. It can be seen that the B-cos explanation highlights important tokens correctly with little focus on irrelevant ones. In contrast, ShapSampl attributes the highest importance to the [SEP] token and provides only little useful information. Meanwhile, DecompX extracts a significant amount of irrelevant information. Overall, we find that the B-cos explanation is more interpretable to humans by providing clearer and more relevant attributions compared to the post-hoc explanations. We provide more examples and analyze two undesired explanations from B-cos LMs in Appendix K.
@@ 4.5 Ablation Study
Table 3: Ablation study of key designs in the B-cos BERT model on HateXplain. Values in parentheses indicate the difference from the full model’s performance. Green (red) indicates the results are better (worse) than the full system.
@@ 5 Comparison of B-cosification Setups
  • Task then B-cos: PLMs are first fine-tuned on a downstream task. B-cos adaptations are then applied, followed by further fine-tuning on the same task for B-cos conversion. This setup is equivalent to Arya et al. (2024) who apply B-cosification to models with downstream task capabilities.
  • B-cos then task: B-cos adaptations are applied to PLMs first, followed by pre-training on unsupervised texts to enhance B-cosification. The pre-trained B-cos models are then fine-tuned on the downstream task.
@@
We evaluate these setups against the B-cosification approach used in our main experiments (B-cos LM) and compare task performance, faithfulness, and training efficiency (cf. Appendix D for B-cos pre-training details). Additionally, we report results for conventional fine-tuning (Conv. LM) and training a randomly initialized B-cos LM (B-cos from scratch). Experiments are conducted on IMDB with B=1.25 for B-cos models, with results averaged over three runs.
Table 4 shows that B-cos LM requires fewer training steps to reach optimal validation performance than conventional fine-tuning. Training B-cos LM from scratch results in worse accuracy and faithfulness, emphasizing the importance of good parameter initialization. Among the two setups that separate task fine-tuning and B-cos conversion, Task then B-cos achieves results similar to B-cos LM but requires more total training steps. B-cos then task initially performs worse under the same training budget. However, with additional pre-training epochs, it surpasses other B-cosification setups in both task performance and faithfulness. Overall, we find that combining task fine-tuning and B-cos conversion is the most efficient approach. However, with sufficient pre-training, B-cos then task can produce more performant and explainable models.
@@ 6 Impact of B-cosification and B Values
Figure 5 shows the effects of varying B on the task performance and explanation faithfulness. Classification accuracy initially improves slightly as B increases from 1 to 1.25, benefiting from the extra non-linearity introduced by B>1. However, beyond this point, accuracy declines as higher alignment pressure reduces model flexibility. A similar trend is observed for Comp, which peaks around B=1.5 before decreasing. This differs from previous findings in vision models (Böhle et al., 2022), which we attribute to the high sparsity of explanations at larger B values. As alignment pressure increases, fewer tokens receive attribution scores that are not close to zero, leading to poor token importance calibration and lower Comp scores. The effects of B on other metrics are similar and can be found in Appendix M.
@@
Since B-cos LMs with larger B values rely on fewer tokens for prediction, we investigate whether this may cause them to overfit and learn biases in the data. For this, we examine label bias and word-level spurious correlations using HateXplain, where approximately 60% of training and test examples have positive labels and societal biases are present. Figure 6 shows that a larger B value (B=2.5) reduces the model capacity, leading to a substantially higher positive rate in predictions and therefore lower class-balanced accuracy. Moreover, the B=2.5 model assigns higher attributions to non-semantic [CLS] and [SEP] tokens, indicating a reduced reliance on meaningful content. Notably, this label bias is not observed in the balanced AG News and IMDB datasets.
We also find that B-cosification, particularly with large B, amplifies reliance on spurious correlations. For example, the prediction positive rate for examples with the word “black” rises from 49.02% in the test set and 52.94% in the conventional model to 59.80%, 56.86%, and 73.53% in B-cos LMs with B=1, 1.5, and, 2.5, respectively (we provide an example in Appendix O). However, the faithfulness and interpretability of B-cos explanations facilitate the detection of spurious correlations and can effectively guide models toward reducing them (Rao et al., 2023). We leave the exploration of B-cos LMs for bias detection and mitigation to future work.
@@ 7 B-cosifying Decoder-Only Models for Generation Tasks
LLMs are increasingly used as general-purpose assistants, with most based on decoder-only architectures (Zhao et al., 2023; Minaee et al., 2024). While our primary focus is on classification tasks using encoder-only models, we also extend B-cosification to decoder-only models for generation tasks to demonstrate the broader applicability of B-cos LMs. Specifically, we apply B-cosification to two decoder-only models, GPT-2 small (Radford et al., 2019, referred to as GPT-2 afterwards) and Llama-3.2-1B (Dubey et al., 2024, referred to as Llama-3.2 afterwards), and evaluate their language modeling performance and explanation quality on two generation tasks. For more details on the datasets, experimental setup and baseline models, see Appendix D.
[…]
B-cosification Setup
Datasets
For explanation evaluation, we use the BLiMP dataset (Warstadt et al., 2020) to assess explanations for linguistic phenomena, and the Indirect Object Identification (IOI) dataset (Brian Muhia, 2022) to test models’ reasoning about object identification. Following Ferrando et al. (2023), we use nine subsets of BLiMP. Each example in both datasets consists of a sentence prefix followed by a target and a foil next-word prediction, differing in whether they align with the phenomenon or ability of interest. Ground truth evidence is provided to support either grammatical correctness or correct object identification. Examples of these datasets can be found in Table 7 in Appendix D.
@@
We evaluate explanation quality using Mean Reciprocal Rank (MRR), where higher scores indicate stronger alignment with the ground truth evidence. To assess language modeling abilities of models, we report two metrics: (1) the probability gap between target and foil predictions, and (2) perplexity (PPL) on a held-out corpus. Following Yin & Neubig (2022), we generate contrastive explanations that explain why the model predicts target tokens instead of foil tokens, and compare B-cos explanations against several baseline methods: L1 gradient norm (Grad Norm), IxG, Occlusion, and two propagation-based methods Logit and ALTI Logit from Ferrando et al. (2023).
@@ 8 Practical Guidance for Training B-cos LMs
Method GPT-2 MRR ($\uparrow$) Llama-3.2 MRR ($\uparrow$)
BLiMP IOI BLiMP IOI
Random 0.5130 0.2360 0.5132 0.2328
Grad Norm 0.5465 0.8599 0.5504 0.3637
IxG 0.4750 0.1112 0.5303 0.1034
Occlusion 0.6365 0.8517 0.6201 0.4767
Logit 0.7307 1.0 - -
ALTI Logit 0.7391 1.0 - -
B-cos 0.7561 1.0 0.6969 0.9913
Table 6: Alignment results (MRR) on BLiMP and IOI. Logit and ALTI Logit results are replicated from the original paper (Ferrando et al., 2023). Best scores are marked in bold. B-cos explanations achieve the best alignment with ground truth evidence. Results for each subset can be found in Table 14 and Table 15 in Appendix P.
@@ 9 Conclusion
In this work, we introduce B-cos LM, a bias-free dynamic linear model that learns task-relevant patterns through increased input-weight alignment pressure. B-cos LMs generate more faithful and human interpretable explanations while maintaining strong task performance and fast convergence. We also explore adapting decoder-only models into B-cos LMs for generation tasks and show that, with additional training, they match the language modeling performance of conventional models while providing better explanations. Based on our in-depth analysis of B-cosification, we also provide a practical guideline for effectively transforming PLMs into B-cos LMs.
@@ 10 Limitations
This study has certain limitations that should be acknowledged. First, the automatic evaluation metrics we use may not fully capture the faithfulness of different explanation methods (Feng et al., 2018; Lapuschkin et al., 2019). However, since there is no universal consensus on the most reliable evaluation metrics, this remains an open challenge in explainability research.
@@
Finally, B-cos explanations, like all input-based approaches, are not suitable for all tasks. They are particularly suitable for NLP tasks where predictions can be directly attributed to specific words or phrases in the input. Examples include various text classification tasks, such as sentiment analysis and toxicity detection, where the presence or absence of certain tokens often provides clear evidence for the model’s decision. In such settings, token-level explanations are both intuitive and precise, enabling users to easily verify whether the model attends to the expected parts of the input. On the contrary, in tasks that involve multi-step reasoning or external knowledge retrieval such as commonsense reasoning or fact verification, token-level explanations are often less informative. In these cases, higher-level explanations, such as concept-based or natural language explanations, may better capture the underlying reasoning process and provide users with more meaningful insights.
@@ Acknowledgements
This work was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – GRK 2853/1 “Neuroexplicit Models of Language, Vision, and Action” - project number 471607914. We also thank all those who provided valuable feedback and suggestions during the development of this work. Their input has been instrumental in improving its quality.
@@ Appendix A Terminology
  • Faithfulness The extent to which an explanation accurately reflects the model’s actual reasoning process (Jacovi & Goldberg, 2020). A faithful explanation should directly correspond to the internal mechanisms that led to the model’s prediction.
  • Human Interpretability The ease with which a person can understand the model’s reasoning from the explanation (Lage et al., 2019). A highly interpretable explanation should be clear, concise, and focused on relevant information while avoiding unnecessary or distracting information. However, an explanation that is easy for humans to interpret may not necessarily reflect the model’s actual reasoning process or align with human reasoning patterns.
@@ Appendix C Dynamic Linear Representation of Model Components
In transformer models, non-linearity is typically introduced using (approximately) piecewise linear activation functions, such as ReLU (Nair & Hinton, 2010) and GELU (Hendrycks & Gimpel, 2016). These functions can be easily interpreted as linear transformations with input-dependent weights. For example, $\text{GELU}(\mathbf{x})=\mathbf{x}\times(0.5+0.5\times\text{erf}(\mathbf{x}/\sqrt{2}))$ can be interpreted as a linear transformation where the second term acts as the dynamic linear weight.
@@
Böhle et al. (2024) showed that attention computations can be seamlessly integrated into B-cos networks as a dynamic linear module:
@@
Following Böhle et al. (2024), we apply bias-free normalization layers to ensure completeness of explanations:

### v4

@@ 4.2 Automatic Evaluation Results
Figure 2: Mean accuracy of conventionally fine-tuned, Saloss and B-cos BERT averaged over three runs. B-cos models perform comparably to conventional models on most tasks.
@@
Table 2 shows the faithfulness scores for post-hoc explanation methods on conventionally fine-tuned and Saloss BERT models, as well as B-cos explanations from B-cos BERT. The results show that B-cos explanations are consistently and substantially more faithful than post-hoc methods across all models and datasets. On average, B-cos explanations outperform the strongest post-hoc methods on conventional models by 14.63 points in Comp and achieve negative Suff scores, indicating that the identified important tokens alone enable even more confident predictions. B-cos also shows significant gains in SeqPG. While Saloss improves faithfulness for some post-hoc methods over conventional models, it still lags markedly behind B-cos LMs. Similar trends are observed for other PLMs (Appendix H) as well. Although we do not include rationale-based models in the main experiments because they typically require additional supervision, a supplementary comparison in Appendix I shows that B-cos BERT still outperforms a rationale-based model on HateXplain.
@@
Table 2: Faithfulness evaluation for conventionally fine-tuned, Saloss, and B-cos BERT across three datasets. The best results are in bold. B-cos explanations are consistently more faithful than post-hoc explanations from both baseline models.
@@ 4.3 Human Evaluation
Contrary to previous B-cos studies that rely solely on automatic evaluations to assess explanations, we conduct the first human study to better evaluate the human interpretability and agreement of B-cos explanations. We compare them against three strong post-hoc explanation methods on conventional BERT models.
Following the practice in enguehard-2023-sequential and dare, we randomly select 50 instances, respectively, from AG News and HateXplain where the B-cos and conventional models make the same prediction. Five annotators then rate the explanations in terms of human interpretability (how well they understand them) and human agreement (how much they agree with them) on a scale of 1-5. Further details on the evaluation criteria, rating scales, annotator instructions, and example annotations can be found in Appendix J.
@@
Figure 3 shows that B-cos explanations have a better human interpretability and exhibit greater alignment with human reasoning than post-hoc methods, even though they are not directly optimized for human agreement. Paired t-tests with a Bonferroni-corrected significance level $\alpha=\frac{0.05}{6}=0.008\overline{3}$ (bonferroni1936teoria) shows that the improvements of B-cos explanations are statistically significant ($p<\alpha$) for both metrics.
@@ 4.4 Qualitative Analysis
Figure 4 provides an example of B-cos and other (post-hoc) explanations. It can be seen that the B-cos explanation highlights important tokens correctly with little focus on irrelevant ones. In contrast, ShapSampl attributes the highest importance to the [SEP] token and provides only little useful information. Meanwhile, DecompX extracts a significant amount of irrelevant information. Overall, the B-cos explanation is more interpretable to humans by providing clearer and more relevant attributions compared to the post-hoc explanations. We provide more examples and analyze two undesired explanations from B-cos LMs in Appendix K.
@@ 4.5 Ablation Study
Table 3: Ablation study of key designs in the B-cos BERT model on HateXplain. Green (red) values in parentheses indicate the results are better (worse) than the full system.
@@ 5 Comparison of B-cosification Setups
  • Task then B-cos: PLMs are first fine-tuned on a downstream task. B-cos adaptations are then applied, followed by further fine-tuning on the same task for B-cos conversion. This setup is equivalent to arya24bcosification who apply B-cosification to models with downstream task capabilities.
  • B-cos then task: B-cos adaptations are applied to PLMs first, followed by unsupervised pre-training to enhance B-cosification. The pre-trained B-cos models are then fine-tuned on the downstream task.
@@
We evaluate these setups against the B-cosification approach used in our main experiments (B-cos LM) and compare task performance, faithfulness, and training efficiency (cf. Appendix D for B-cos pre-training details). We also report results for conventional fine-tuning (Conv. LM) and randomly initialized B-cos models (B-cos from scratch). All experiments are run on IMDB with B=1.25 for B-cos models, with results averaged over three runs.
Table 4 shows that B-cos LM requires fewer training steps to reach optimal validation performance than conventional fine-tuning. Training B-cos LM from scratch results in worse accuracy and faithfulness, showing the importance of good parameter initialization. Among the two setups that separate task fine-tuning and B-cos conversion, Task then B-cos achieves results similar to B-cos LM but requires more total training steps. B-cos then task initially performs worse under the same training budget. However, with additional pre-training epochs, it surpasses other B-cosification setups in both task performance and faithfulness. Overall, we find that combining task fine-tuning and B-cos conversion is the most efficient approach. However, with sufficient pre-training, B-cos then task can produce more performant and explainable models.
@@ 6 Impact of B-cosification and B Values
Figure 5 shows the effects of varying B on the task performance and explanation faithfulness. Classification accuracy initially improves slightly as B increases from 1 to 1.25, benefiting from the extra non-linearity introduced by B>1. However, beyond this point, accuracy declines as higher alignment pressure reduces model flexibility. A similar trend is observed for Comp, which peaks around B=1.5 before decreasing. This differs from previous findings in vision models (Bohle_2022_CVPR), which we attribute to the high sparsity of explanations at larger B values. As alignment pressure increases, fewer tokens receive attribution scores that are not close to zero, leading to poor token importance calibration and worse Comp scores. The effects of B on other metrics are similar and can be found in Appendix M.
@@
PLMs often acquire biased patterns present in their training data (wang-demberg-2024-parameter; wang-demberg-2024-rsa). Since B-cos LMs with larger B values rely on fewer tokens for prediction, we investigate whether this may cause them to overfit and learn biases in the data. For this, we examine label bias and word-level spurious correlations using HateXplain, where approximately 60% of training and test examples have positive labels and societal biases are present. Figure 6 shows that a larger B value (B=2.5) reduces the model capacity, leading to a substantially higher positive rate in predictions and therefore lower class-balanced accuracy. Moreover, the B=2.5 model assigns higher attributions to non-semantic [CLS] and [SEP] tokens, indicating a reduced reliance on meaningful content. Notably, this label bias is not observed in the balanced AG News and IMDB datasets.
[…]
We also find that B-cosification, particularly with large B, amplifies reliance on spurious correlations. For example, the prediction positive rate for examples with the word “black” rises from 49.02% in the test set and 52.94% in the conventional model to 59.80%, 56.86%, and 73.53% in B-cos LMs with B=1, 1.5, and, 2.5, respectively (we provide an example in Appendix O). However, the faithfulness and interpretability of B-cos explanations facilitate the detection of spurious correlations and can effectively guide models toward reducing them (Rao_2023_ICCV; balkir-etal-2022-challenges; bridge-wang-2025). We leave the exploration of B-cos LMs for bias detection and mitigation to future work.
@@ 7 B-cosifying Decoder-Only Models for Generation Tasks
LLMs are increasingly used as general-purpose assistants, with most based on decoder-only architectures (llm-survey-2023; llm-survey-2024). While our primary focus is on classification tasks using encoder-only models, we also extend B-cosification to decoder-only models for generation tasks to demonstrate the broader applicability of B-cos LMs. Specifically, we apply B-cosification to two decoder-only models, GPT-2 small (Radford-2019-GPT, referred to as GPT-2 afterwards) and Llama-3.2-1B (Llama3, referred to as Llama-3.2 afterwards), and evaluate their language modeling performance and explanation quality on two generation tasks. For more details on the datasets, experimental setup and baseline models, see Appendix D.
@@
B-cosification Setup
Datasets
For explanation evaluation, we use the BLiMP dataset (warstadt2020blimp) to assess explanations for linguistic phenomena, and the Indirect Object Identification (IOI) dataset (fahamu_2022) to test models’ reasoning about object identification. Following ferrando-etal-2023-explaining, we use nine subsets of BLiMP. Each example in both datasets consists of a sentence prefix followed by a target and a foil next word prediction, differing in whether they align with the phenomenon or ability of interest. Ground truth evidence is provided to support either grammatical correctness or correct object identification. Examples of these datasets can be found in Table 7 in Appendix D.
Method GPT-2 MRR ($\uparrow$) Llama-3.2 MRR ($\uparrow$)
BLiMP IOI BLiMP IOI
Random 0.5130 0.2360 0.5132 0.2328
Grad Norm 0.5465 0.8599 0.5504 0.3637
IxG 0.4750 0.1112 0.5303 0.1034
Occlusion 0.6365 0.8517 0.6201 0.4767
Logit 0.7307 1.0 - -
ALTI Logit 0.7391 1.0 - -
B-cos 0.7561 1.0 0.6969 0.9913
Table 6: Alignment results (MRR) on BLiMP and IOI. Logit and ALTI Logit results are replicated from the original paper (ferrando-etal-2023-explaining). Best scores are marked in bold. B-cos explanations achieve the best alignment with ground truth evidence. Results for each subset can be found in Table 14 and Table 15 in Appendix P.
@@
We evaluate explanation quality using Mean Reciprocal Rank (MRR), where higher scores indicate stronger alignment with the ground truth evidence. To assess language modeling abilities of models, we report two metrics: (1) the probability gap between target and foil predictions, and (2) perplexity (PPL) on a held-out corpus. Following yin-neubig-2022-interpreting, we generate contrastive explanations that explain why the model predicts target tokens instead of foil tokens, and compare B-cos explanations against several baseline methods: L1 gradient norm (Grad Norm), IxG, Occlusion, and two propagation-based methods Logit and ALTI Logit from ferrando-etal-2023-explaining.
@@ 9 Conclusion
In this work, we introduce B-cos LM, a bias-free dynamic linear model that learns task-relevant patterns through increased input-weight alignment pressure. B-cos LMs generate more faithful and human interpretable explanations while maintaining strong task performance and fast convergence. We further explore adapting decoder-only models into B-cos LMs for generation tasks and show that, with additional training, they match the language modeling performance of conventional models while providing better explanations. Finally, based on our systematic analysis, we provide practical guidelines for effectively transforming PLMs into B-cos LMs.
@@ 10 Limitations
This study has certain limitations that should be acknowledged. First, the automatic evaluation metrics we use may not fully capture the faithfulness of different explanation methods (feng-etal-2018-pathologies; lapuschkin2019unmasking). However, since there is no universal consensus on the most reliable evaluation metrics, this remains an open challenge in explainability research.
@@
Finally, B-cos explanations, like all input-based approaches, are not suitable for all tasks. They are particularly suitable for NLP tasks where predictions can be directly attributed to specific words or phrases. Examples include various text classification tasks, such as sentiment analysis and toxicity detection, where the presence or absence of certain tokens often provides clear evidence for the model’s decision. Conversely, for tasks requiring multi-step reasoning or external knowledge (e.g., commonsense reasoning, fact verification), token-level rationales become less informative. Higher-level explanations, such as concept-based or natural language explanations, may instead offer more meaningful insight into the model’s reasoning.
@@ Acknowledgements
This work was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – GRK 2853/1 "Neuroexplicit Models of Language, Vision, and Action" - project number 471607914. We also thank all those who provided valuable feedback and suggestions during the development of this work. Their input has been instrumental in improving its quality.
@@ Appendix A Terminology
  • Faithfulness The extent to which an explanation accurately reflects the model’s actual reasoning process (jacovi-goldberg-2020-towards). A faithful explanation should directly correspond to the internal mechanisms that led to the model’s prediction.
  • Human Interpretability The ease with which a person can understand the model’s reasoning from the explanation (DBLP:journals/corr/abs-1902-00006). A highly interpretable explanation should be clear, concise, and focused on relevant information while avoiding unnecessary or distracting information. However, an explanation that is easy for humans to interpret may not necessarily reflect the model’s actual reasoning process or align with human reasoning patterns.
@@ Appendix C Dynamic Linear Representation of Model Components
In transformer models, non-linearity is typically introduced using (approximately) piecewise linear activation functions, such as ReLU (DBLP:conf/icml/NairH10) and GELU (DBLP:journals/corr/HendrycksG16). These functions can be easily interpreted as linear transformations with input-dependent weights. For example, $\text{GELU}(\mathbf{x})=\mathbf{x}\times(0.5+0.5\times\text{erf}(\mathbf{x}/\sqrt{2}))$ can be interpreted as a linear transformation where the second term acts as the dynamic linear weight.
@@
bohle2024b showed that attention computations can be seamlessly integrated into B-cos networks as a dynamic linear module:
@@
Following bohle2024b, we apply bias-free normalization layers to ensure completeness of explanations: